OpenStudy (anonymous):

square root of 49x^16y^14

OpenStudy (anonymous):

sqrt(49) = 7 sqrt(x^16) = x^8 (becuase x^8 times x^8 = x^16) sqrt (y^14) = y^7 (becuase y^7 times y^7 = y^14) so you get 7x^8y^7.